# McKean-Vlasov equations and nonlinear Fokker-Planck equations with critical singular Lorentz kernels

### v1

@@ Abstract.
We prove the existence and conditional uniqueness in the Krylov class for SDEs with singular divergence-free drifts in the endpoint critical Lorentz space $L^{\infty}(0,T;L^{d,\infty}({\mathbb{R}}^{d}))$, $d\geqslant 2$, which particularly includes the Biot-Savart law. The uniqueness result is shown to be optimal in dimensions $d\geqslant 3$, by constructing different martingale solutions in the case of supercritical Lorentz drifts. As a consequence, the well-posedness of McKean-Vlasov equations and nonlinear Fokker-Planck equations with critical singular kernels is derived. In particular, this yields the uniqueness of the 2D vorticity Navier-Stokes equations even in certain supercritical-scaling spaces. Furthermore, we prove that the path laws of solutions to McKean–Vlasov equations with critical singular kernel form a nonlinear Markov process in the sense of McKean.
@@ 1. Introduction and main results
[…]
with $K(t,x)=K_{\mathrm{BS}}(x)$ given by (1.1). The marginal law of McKean-Vlasov solutions satisfies the $2$D vorticity (NSE), and thus, (1.1)-(MVE) provide a probabilistic interpretation of (NSE). We note that, rather than in the usual space $L^{2}_{loc}({\mathbb{R}}^{2})$, the Biot-Savart law belongs to the critical Lorentz space $L^{2,\infty}({\mathbb{R}}^{2})$. Hence, the nonlinearity in (NSE) involves a singular kernel. Therefore, achieving the McKean picture for the 2D vorticity NSE, or more general nonlinear Fokker-Planck equations
@@
where the interaction kernel $K$ can be singular and time-dependent, requires the solvability theory for McKean-Vlasov equations (MVE) with singular kernels. Although the well-posedness of the $2$D vorticity NSE has been extensively studied (see, e.g., [16, 18, 19, 21, 24, 5]), the singularity and criticality of kernels pose a major challenge to solve singular McKean-Vlasov equations. As a matter of fact, very few results are known in the case of critical singular kernels in dimensions $d\geqslant 2$.
@@
The remaining $2$D case is more subtle than the higher dimensional case. Actually, even in the simple case of SDEs with time-independent drifts, the standard weak sector condition is unclear for the corresponding Dirichlet form in dimension two, while it is valid in dimensions $d\geqslant 3$. Thus, solving singular SDEs with the endpoint critical $L^{\infty}_{t}L^{2,\infty}_{x}$-Lorentz drifts in dimension two, closely related to the Biot-Savart law, remains a challenging open problem.
@@
    The uniqueness is proved for the 2D vorticity NSE in certain Krylov class, which includes a supercritical scaling space $L^{\mathfrak{q}^{\prime}}_{t}L^{\mathfrak{p}^{\prime}}_{x}$ with $11/\mathfrak{p}^{\prime}+1/\mathfrak{q}^{\prime}>$ 1.
@@
    The nonlinear Markov property of solutions is derived in the Krylov class $L^{\mathfrak{q}^{\prime}}_{t}L^{\mathfrak{p}^{\prime}}_{x}$ with $11/\mathfrak{p}^{\prime}+1/\mathfrak{q}^{\prime}>$ 1 for (MVE) associated to the 2D vorticity NSE.
@@ 1.1.1. SDEs with endpoint critical drifts
:=${\mathscr{I}}:=\left\{(p,q)\in(1,\infty)^{2}:\frac{d}{p}+\frac{2}{q}<2\right\}.$
@@ 1.1.2. Optimality of well-posedness
    The uniqueness result in Theorem 1.1 is important to establish the uniqueness of solutions to McKean-Vlasov equations and nonlinear Fokker-Planck equations with endpoint critical kernels. In particular, the Krylov class in Theorem 1.1 enables to derive the uniqueness of solutions to the $2$D vorticity NSE in certain supercritical spaces. See Theorem 1.8 below.
@@ 1.1.3. McKean-Vlasov equations and nonlinear Fokker-Planck equations with critical kernels
    There exist $(\mathfrak{p},\mathfrak{q})\in{\mathscr{I}}$ and $\varepsilon_{0}>0$, depending on $d$ and $\|K\|_{L^{\infty}_{t}L_{x}^{d,\infty}}$, such that if $\zeta_{a}\in\mathcal{M}(\mathbb{R}^{d})$ with $|\zeta_{a}|({\mathbb{R}}^{d})\leqslant\varepsilon_{0}$, then the narrowly continuous weak solution to equation (NFPE) is unique in the class of functions $\rho$ satisfying the following property: for each $\delta\in(0,T)$,
    $\rho\in L^{\mathfrak{q}^{\prime}}(\delta,T;L^{\mathfrak{p}^{\prime}}_{x}),$
    where $(\mathfrak{p}^{\prime},\mathfrak{q}^{\prime})=(\mathfrak{p}/(\mathfrak{p}-1),\mathfrak{q}/(\mathfrak{q}-1))$.
@@
where $\mathfrak{p}^{\prime}$ and $\mathfrak{q}^{\prime}$ are the same exponents as in Theorem 1.7 satisfying $1{1}/{\mathfrak{p}^{\prime}}+{1}/{\mathfrak{q}^{\prime}}>$ 1.
@@
    Actually, the $2$D vorticity NSE has the invariant scaling
@@
    Note that, the Biot-Savart kernel is invariant under the scaling, namely, $K_{\lambda}(t,x)=K(t,x)$, $\forall\lambda>0$. The above scaling leaves the mixed Lebesgue space $L^{q}(\mathbb{R}_{+};L^{p}_{x})$ invariant if $1{1}/{p}+{1}/{q}=$ 1. For the exponents $(\mathfrak{p}^{\prime},\mathfrak{q}^{\prime})$ in (1.4), since $1{1}/{\mathfrak{p}^{\prime}}+{1}/{\mathfrak{q}^{\prime}}>$ 1, one has
@@ 1.3. Ideas of the proof
for certain exponents $l>2$ and $1<q<2l/(l-2)$. It is important that the spatial integrability exponent here can be raised above $2$.
@@
where the exponents $\mathfrak{l}>2$ and $\mathfrak{q}>2\mathfrak{l}/(\mathfrak{l}-2)$. In both steps, it is quite delicate to select appropriate integrability exponents to ensure that the final upgraded integrability exponents $(\mathfrak{p},\mathfrak{q})\in\mathscr{I}$, i.e., $1{1}/{\mathfrak{p}}+{1}/{\mathfrak{q}}<$ 1, as required by the Krylov class.
@@
More precisely, we construct a divergence-free vector field $b\in L_{x}^{p,\infty}$, $p\in(d/2,d)$. It is anti-symmetric with respect to the hyperplane $\Pi=\{x\in{\mathbb{R}}^{d}:x_{d}=0\}$, and when $x$ is located in the cone ${\mathcal{C}}_{\kappa}=\{x:x_{d}>\kappa|\widehat{x}|\}$ with $\kappa\geqslant 1$, it can be very singular in the direction of the last coordinate
@@
Then, we consider two sequences $(x^{n})\subseteq\{\hat{x}=0\}$ and $(y^{n})\subseteq\Pi$ converging to the singular point 00 in different ways. Because of the anti-symmetry of the drift, the corresponding martingale solutions $\{\mathbb{P}_{y^{n}}\}$ are invariant under reflection across the hyperplane $\Pi$, and so is the limit $\mathbb{P}_{0}$.
In contrast, for the martingale solution $\mathbb{P}_{x^{n}}$, the singularity of the drift along the axis $x_{d}$ forces the solution trajectories starting from the small cone ${\mathcal{C}}_{\kappa}$, with a uniform probability (independent of $n$), to stay in the larger cone ${\mathcal{C}}_{1}$ and return back to ${\mathcal{C}}_{\kappa}$ after certain time. As a result, the trajectories are more likely to intersect with the hyperplane $\{x_{d}=2\}$ before exiting ${\mathcal{C}}_{1}$, and to remain in the region $\{x_{d}>1\}$ for a unit time. Intuitively, as illustrated in Figure 2, the solution process follows the green trajectories with high probability. Consequently, the limit ~$\widetilde{\mathbb{P}}_{0}$ of $(\mathbb{P}_{x^{n}})$ is more concentrated on the half space $\Pi_{+}=\{x\in{\mathbb{R}}^{d}:x_{d}>0\}$, and thus, leads to the non-uniqueness ~$\widetilde{\mathbb{P}}_{0}\not=\mathbb{P}_{0}$.
@@
Let $T$ be any fixed time horizon. For $p\in[1,\infty]$, $p^{\prime}$ denotes its conjugate number $\frac{p}{p-1}$. For $z=(t,x)\in{\mathbb{R}}\times{\mathbb{R}}^{d}$, we set $Q_{R}(z):=(t-R^{2},t)\times B_{R}(x)$ and $Q_{R}:=(-R^{2},0)\times B_{R}$. For any $p,q\in(1,\infty)$, let
@@ 2.1. Aronson-type estimate
Suppose that $w$ is a nonnegative, nondecreasing continuous function on $[0,\infty)$. Let $p\geqslant 2,a,\beta,\Gamma>0$ and $\delta\in(0,1]$. Then, there exists $C$ depending only on $a$ and $\beta$ such that if
@@
[…]
@@
[…]
@@
[…]
@@ 2.2. Hölder regularity estimate
where we also used the fact that $-\frac{d}{2}(\frac{1}{p}-\frac{1}{r})(1-\frac{1}{q})^{-1}>-1$ in the last step, due to the conditions on $p$, $q$ and $r$. Thus, (2.15) is proved. ∎
@@ 2.3. Sobolev regularity estimate
Let $p,q\in(1,\infty)$, $r\in[p,\infty)$ and $s\in[q,\infty)$. Assume $\partial_{t}u\in L^{q}_{t}L_{x}^{p}$, $u\in L^{q}_{t}W^{2,p}_{x}$. If $11<2/q+d/p=2/s+d/r+$ 1, then
@@
$f\in L_{t}^{\infty}(L_{x}^{1}\cap L_{x}^{d})$$u\in L_{t}^{\infty}L_{x}^{r,1}$$r\gg\frac{d}{d-2}$$bu\in L_{t}^{q}L_{x}^{l,1}$$l=\frac{dr}{d+r},q\gg 1$$\nabla u\in L_{t}^{q}L_{x}^{l}$$b\cdot\nabla u\in L_{t}^{q}L_{x}^{p}$$p=\frac{dl}{d+l}$$\nabla^{2}u\in L_{t}^{q}L_{x}^{p}$$\nabla u\in L_{t}^{\mathfrak{q}}L_{x}^{\mathfrak{l},1}$$\delta\in(0,\frac{1}{2}),\mathfrak{l}=\frac{p}{\delta},\mathfrak{q}=\frac{q}{\delta}$$b\cdot\nabla u\in L_{t}^{\mathfrak{q}}L_{x}^{\mathfrak{p}}$$\mathfrak{p}=\frac{d\mathfrak{l}}{d+\mathfrak{l}}$$u\in L_{t}^{\mathfrak{q}}W_{x}^{2,\mathfrak{p}},\quad\partial_{t}u\in L_{t}^{\mathfrak{q}}L_{x}^{\mathfrak{p}}$rop.2.6Prop. 2.8Prop. 2.8Prop. B.2Prop. 2.6Prop. 2.8
@@
$\|{{-}}f\|{{-}}_{L^{p}(A)}:=\left(\fint_{A}|f|^{p}\right)^{\frac{1}{p}}=\left(\frac{1}{|A|}\int_{A}|f|^{p}\right)^{\frac{1}{p}}.$
@@
$\|{{-}}v\|{{-}}_{L^{p}({Q_{R/2}(z))}}\leqslant\theta\|{{-}}v\|{{-}}_{L^{p}(Q_{R}(z))}+C_{2}\|{{-}}v\|{{-}}_{L^{1}(Q_{R}(z))}+C_{2}\|{{-}}g\|{{-}}_{L^{p}(Q_{R}(z))}$
@@
$\|{{-}}v\|{{-}}_{L^{q}(Q_{R/2}(z))}\leqslant C\|{{-}}v\|{{-}}_{L^{p}({Q_{R}(z))}}+C\|{{-}}g\|{{-}}_{L^{q}(Q_{R}(z))},$ (2.23)
@@
$-\iint_{Q_{R}}\Delta u~{}\bar{u}\varphi^{10}\phi^{2}=\iint_{Q_{R}}|\nabla u|^{2}\varphi^{10}\phi^{2}+10\iint_{Q_{R}}\bar{u}\varphi^{9}\phi^{2}~{}\nabla u\cdot\nabla\varphi.$
@@
[…]
@@
[…]
@@
[…]
@@
Let $d=2$. There exists a universal constant $\gamma>2$ only depending on $\|b\|_{L^{\infty}_{t}L^{2,\infty}_{x}}$ such that for any weak solution to (2.11) on $[0,T]\times{\mathbb{R}}^{d}$ with $c=0$, it holds
@@
$\|{{-}}v\|{{-}}_{L^{p}(Q_{{R}/{2}})}\leqslant\frac{1}{2}\|{{-}}v\|{{-}}_{L^{p}(Q_{R})}+C_{6}\|{{-}}v\|{{-}}_{L^{1}(Q_{R})}+C_{6}\|{{-}}g\|{{-}}_{L^{p}(Q_{R})},$
@@
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}:=\|f\|_{L^{\infty}_{t}L_{x}^{1}}+\|f\|_{L^{\infty}_{t}L_{x}^{2}}.$
@@
$\|\nabla{\mathcal{T}}(f)\|_{L^{\nu}_{t}L_{x}^{\mu}}\leqslant C\left(\|\partial_{t}{\mathcal{T}}(f)\|_{L^{\eta}_{t,x}}+\|\nabla^{2}{\mathcal{T}}(f)\|_{L^{\eta}_{t,x}}\right)\leqslant C\|f\|_{L^{\eta}_{t,x}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$
@@
$\|u\|_{L^{\infty}_{t}L_{x}^{r,1}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\quad\text{with}\quad r=\left(\frac{1}{\mu}-\frac{1}{2}\right)^{-1}.$
@@
$\|bu\|_{L^{\nu}_{t}L_{x}^{\mu}}\leqslant C\|b\|_{L^{\infty}_{t}L_{x}^{2,\infty}}\|u\|_{L^{\infty}_{t}L_{x}^{r,1}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$ (2.32)
@@
$\|\nabla u\|_{L^{\nu}_{t}L^{\mu}_{x}}\leqslant C\|\nabla^{2}{\mathcal{T}}(bu)\|_{L^{\nu}_{t}L^{\mu}_{x}}+C\|\nabla{\mathcal{T}}(f)\|_{L^{\nu}_{t}L^{\mu}_{x}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$ (2.33)
@@
$\|\nabla u\|_{L^{q}_{t}L^{l,1}_{x}}\leqslant\|\nabla u\|_{L^{\nu}_{t}L^{\mu}_{x}}^{\theta}\|\nabla u\|_{L^{\gamma}_{t,x}}^{1-\theta}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|},$
@@
$\|b\cdot\nabla u\|_{L^{q}_{t}L_{x}^{p}}\leqslant C\|b\|_{L^{\infty}_{t}L_{x}^{2,\infty}}\|\nabla u\|_{L^{q}_{t}L_{x}^{l,1}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$ (2.35)
@@
$\|\nabla^{2}u\|_{L^{q}_{t}L_{x}^{p}}\leqslant C\left(\|b\cdot\nabla u\|_{L^{q}_{t}L_{x}^{p}}+\|f\|_{L^{q}_{t}L_{x}^{p}}\right)\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$
@@
$\|\nabla u\|_{L^{\mathfrak{q_{i}}}_{t}L_{x}^{\mathfrak{l}_{i}}}\leqslant\|\nabla^{2}u\|_{L^{q}_{t}L_{x}^{p}}^{\delta_{i}}\|u\|_{L^{\infty}_{t}C^{\alpha_{i}}}^{1-\delta_{i}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|},$

### v2

@@ Abstract.
We prove the existence and conditional uniqueness in the Krylov class for SDEs with singular divergence-free drifts in the endpoint critical Lorentz space $L^{\infty}(0,T;L^{d,\infty}(\mathbb{R}^{d}))$, $d\geqslant 2$, which particularly includes the 22D Biot-Savart law. The uniqueness result is shown to be optimal in dimensions $d\geqslant 3$, by constructing different martingale solutions in the case of supercritical Lorentz drifts. As a consequence, the well-posedness of McKean-Vlasov equations and nonlinear Fokker-Planck equations with critical singular kernels is derived. In particular, this yields the uniqueness of the 2D vorticity Navier-Stokes equations even in certain supercritical-scaling spaces. Furthermore, we prove that the path laws of solutions to McKean–Vlasov equations with critical singular kernel form a nonlinear Markov process in the sense of McKean.
@@ 1. Introduction and main results
$\left\{\begin{aligned} &\mathrm{d}X_{t}=\int_{\mathbb{R}^{d}}K(t,X_{t}-y)\rho(t,\mathrm{d}y)\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\\ &\rho(t)=\mathrm{law}(X_{t}),\quad\rho|_{t=0}=\zeta\in\mathcal{P}(\mathbb{R}^{d})\end{aligned}\right$. (MVE)
with $K(t,x)=K_{\mathrm{BS}}(x)$ given by (1.1). The marginal law of McKean-Vlasov solutions satisfies the 22D vorticity (NSE), and thus, (1.1)-(MVE) provide a probabilistic interpretation of (NSE). We note that, rather than in the usual space $L^{2}_{loc}({\mathbb{R}}^{2})$, the Biot-Savart law belongs to the critical Lorentz space $L^{2,\infty}({\mathbb{R}}^{2})$. Hence, the nonlinearity in (NSE) involves a singular kernel. Therefore, achieving the McKean picture for the 2D vorticity NSE, or more general nonlinear Fokker-Planck equations
@@
where the interaction kernel $K$ can be singular and time-dependent, requires the solvability theory for McKean-Vlasov equations (MVE) with singular kernels. Although the well-posedness of the 22D vorticity NSE has been extensively studied (see, e.g., [16, 18, 19, 21, 24, 5]), the singularity and criticality of kernels pose a major challenge to solve singular McKean-Vlasov equations. As a matter of fact, very few results are known in the case of critical singular kernels in dimensions $d\geqslant 2$.
@@
The remaining 22D case is more subtle than the higher dimensional case. Actually, even in the simple case of SDEs with time-independent drifts, the standard weak sector condition is unclear for the corresponding Dirichlet form in dimension two, while it is valid in dimensions $d\geqslant 3$. Thus, solving singular SDEs with the endpoint critical $L^{\infty}_{t}L^{2,\infty}_{x}-$Lorentz drifts in dimension two, closely related to the Biot-Savart law, remains a challenging open problem.
@@
    The uniqueness is proved for the 2D vorticity NSE in certain Krylov class, which includes a supercritical scaling space $L^{\mathfrak{q}^{\prime}}_{t}L^{\mathfrak{p}^{\prime}}_{x}$ with $1/\mathfrak{p}^{\prime}+1/\mathfrak{q}^{\prime}>1$.
@@
    The nonlinear Markov property of solutions is derived in the Krylov class $L^{\mathfrak{q}^{\prime}}_{t}L^{\mathfrak{p}^{\prime}}_{x}$ with $1/\mathfrak{p}^{\prime}+1/\mathfrak{q}^{\prime}>1$ for (MVE) associated to the 2D vorticity NSE.
@@ 1.1.1. SDEs with endpoint critical drifts
${\mathscr{I}}:=\left\{(p,q)\in(1,\infty)^{2}:\frac{d}{p}+\frac{2}{q}<2\right\}.$
@@ 1.1.2. Optimality of well-posedness
    The uniqueness result in Theorem 1.1 is important to establish the uniqueness of solutions to McKean-Vlasov equations and nonlinear Fokker-Planck equations with endpoint critical kernels. In particular, the Krylov class in Theorem 1.1 enables to derive the uniqueness of solutions to the 22D vorticity NSE in certain supercritical spaces. See Theorem 1.8 below.
@@ 1.1.3. McKean-Vlasov equations and nonlinear Fokker-Planck equations with critical kernels
    There exist $(\mathfrak{p},\mathfrak{q})\in{\mathscr{I}}$ and $\varepsilon_{0}>0$, depending on $d$ and $\|K\|_{L^{\infty}_{t}L_{x}^{d,\infty}}$, such that if $\zeta_{a}\in\mathcal{M}(\mathbb{R}^{d})$ with $|\zeta_{a}|({\mathbb{R}}^{d})\leqslant\varepsilon_{0}$, then the narrowly continuous weak solution to equation (NFPE) is unique in the class of functions $\rho$ satisfying the following property: for each $\delta\in(0,T)$, ρL^q’(δ, T; L^p’_x), where $(\mathfrak{p}^{\prime},\mathfrak{q}^{\prime})=(\mathfrak{p}/(\mathfrak{p}-1),\mathfrak{q}/(\mathfrak{q}-1))$.
@@
where $\mathfrak{p}^{\prime}$ and $\mathfrak{q}^{\prime}$ are the same exponents as in Theorem 1.7 satisfying ${1}/{\mathfrak{p}^{\prime}}+{1}/{\mathfrak{q}^{\prime}}>1$.
@@
    Actually, the 22D vorticity NSE has the invariant scaling
@@
    Note that, the Biot-Savart kernel is invariant under the scaling, namely, $K_{\lambda}(t,x)=K(t,x)$, $\forall\lambda>0$. The above scaling leaves the mixed Lebesgue space $L^{q}(\mathbb{R}_{+};L^{p}_{x})$ invariant if 1/p+1/q=1{1}/{p}+{1}/{q}=1. For the exponents $(\mathfrak{p}^{\prime},\mathfrak{q}^{\prime})$ in (1.4), since ${1}/{\mathfrak{p}^{\prime}}+{1}/{\mathfrak{q}^{\prime}}>1$, one has
@@ 1.3. Ideas of the proof
for certain exponents $l>2$ and 1<q<2l/(l2)1<q<2l/(l-2). It is important that the spatial integrability exponent here can be raised above $2$.
@@
where the exponents $\mathfrak{l}>2$ and $\mathfrak{q}>2\mathfrak{l}/(\mathfrak{l}-2)$. In both steps, it is quite delicate to select appropriate integrability exponents to ensure that the final upgraded integrability exponents $(\mathfrak{p},\mathfrak{q})\in\mathscr{I}$, i.e., ${1}/{\mathfrak{p}}+{1}/{\mathfrak{q}}<1$, as required by the Krylov class.
@@
[…]
@@
Then, we consider two sequences $(x^{n})\subseteq\{\hat{x}=0\}$ and $(y^{n})\subseteq\Pi$ converging to the singular point 0 in different ways. Because of the anti-symmetry of the drift, the corresponding martingale solutions $\{\mathbb{P}_{y^{n}}\}$ are invariant under reflection across the hyperplane $\Pi$, and so is the limit $\mathbb{P}_{0}$.
In contrast, for the martingale solution $\mathbb{P}_{x^{n}}$, the singularity of the drift along the axis $x_{d}$ forces the solution trajectories starting from the small cone ${\mathcal{C}}_{\kappa}$, with a uniform probability (independent of $n$), to stay in the larger cone ${\mathcal{C}}_{1}$ and return back to ${\mathcal{C}}_{\kappa}$ after certain time. As a result, the trajectories are more likely to intersect with the hyperplane $\{x_{d}=2\}$ before exiting ${\mathcal{C}}_{1}$, and to remain in the region $\{x_{d}>1\}$ for a unit time. Intuitively, as illustrated in Figure 2, the solution process follows the green trajectories with high probability. Consequently, the limit $\widetilde{\mathbb{P}}_{0}$ of $(\mathbb{P}_{x^{n}})$ is more concentrated on the half space $\Pi_{+}=\{x\in{\mathbb{R}}^{d}:x_{d}>0\}$, and thus, leads to the non-uniqueness $\widetilde{\mathbb{P}}_{0}\not=\mathbb{P}_{0}$.
@@
Let $T$ be any fixed time horizon. For $\in[1,\infty], pp^{\prime}$ denotes its conjugate number $\frac{p}{p-1}$. For $z=(t,x)\in{\mathbb{R}}\times{\mathbb{R}}^{d}$, we set $Q_{R}(z):=(t-R^{2},t)\times B_{R}(x)$ and $Q_{R}:=(-R^{2},0)\times B_{R}$. For any $p,q\in(1,\infty)$, let
@@ 2.1. Aronson-type estimate
Suppose that $w$ is a nonnegative, nondecreasing continuous function on $0,\infty)$. Let $p\geqslant 2,a,\beta,\Gamma>0$ and $\delta\in(0,1]$. Then, there exists $C$ depending only on $a$ and $\beta$ such that if
@@
${\leqslant}$ $-\left(4-2/p\right)\|\nabla f_{t}^{p}\|_{L^{2}_{x}}^{2}+p|\alpha|^{2}\|f_{t}^{p}\|_{L^{2}_{x}}^{2}+Cp|\alpha|\|b\|_{L_{t}^{\infty}L_{x}^{d,\infty}}\|f_{t}^{p}\|_{L^{2}_{x}}\|\nabla f_{t}^{p}\|_{L^{2}_{x}}$
@@
${\leqslant}$ $-c\frac{\|f_{t}^{p}\|_{L^{2}_{x}}^{2+\frac{4}{d}}}{\|f_{t}^{p}\|_{L^{1}_{x}}^{\frac{d}{4}}}+\left(1+C\|b\|^{2}_{L_{t}^{\infty}L_{x}^{d,\infty}}\right)p^{2}|\alpha|^{2}\|f_{t}^{p}\|_{L^{2}_{x}}^{2},$
@@
${\leqslant}$ $C\|f\mu^{\frac{2}{3}}\|_{L_{x}^{2}}^{\frac{1}{2}}\|\nabla(f\mu^{\frac{2}{3}})\|_{L_{x}^{2}}^{\frac{1}{2}}$
@@ 2.2. Hölder regularity estimate
where we also used the fact that $\frac{d}{2}(\frac{1}{p}-\frac{1}{r})(1-\frac{1}{q})^{-1}>-1$ in the last step, due to the conditions on $p$, $q$ and $r$. Thus, (2.15) is proved. ∎
@@ 2.3. Sobolev regularity estimate
Let $p,q\in(1,\infty)$, $r\in[p,\infty)$ and $s\in[q,\infty)$. Assume $\partial_{t}u\in L^{q}_{t}L_{x}^{p}$, $u\in L^{q}_{t}W^{2,p}_{x}$. If $1<2/q+d/p=2/s+d/r+1$, then
@@
$f\in L_{t}^{\infty}(L_{x}^{1}\cap L_{x}^{d})$$u\in L_{t}^{\infty}L_{x}^{r,1}$$r\gg\frac{d}{d-2}$$bu\in L_{t}^{q}L_{x}^{l,1}$$l=\frac{dr}{d+r},q\gg 1$$\nabla u\in L_{t}^{q}L_{x}^{l}$$b\cdot\nabla u\in L_{t}^{q}L_{x}^{p}$$p=\frac{dl}{d+l}$$\nabla^{2}u\in L_{t}^{q}L_{x}^{p}$$\nabla u\in L_{t}^{\mathfrak{q}}L_{x}^{\mathfrak{l},1}$$\delta\in(0,\frac{1}{2}),\mathfrak{l}=\frac{p}{\delta},\mathfrak{q}=\frac{q}{\delta}$$b\cdot\nabla u\in L_{t}^{\mathfrak{q}}L_{x}^{\mathfrak{p}}$$\mathfrak{p}=\frac{d\mathfrak{l}}{d+\mathfrak{l}}$$u\in L_{t}^{\mathfrak{q}}W_{x}^{2,\mathfrak{p}},\quad\partial_{t}u\in L_{t}^{\mathfrak{q}}L_{x}^{\mathfrak{p}}$Prop.2.6Prop. 2.8Prop. 2.8Prop. B.2Prop. 2.6Prop. 2.8
@@
$\|\hskip-6.4pt{{-}}f\|\hskip-6.4pt{{-}}_{L^{p}(A)}:=\left(\fint_{A}|f|^{p}\right)^{\frac{1}{p}}=\left(\frac{1}{|A|}\int_{A}|f|^{p}\right)^{\frac{1}{p}}.$
@@
$\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{p}({Q_{R/2}(z))}}\leqslant\theta\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{p}(Q_{R}(z))}+C_{2}\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{1}(Q_{R}(z))}+C_{2}\|\hskip-6.4pt{{-}}g\|\hskip-6.4pt{{-}}_{L^{p}(Q_{R}(z))}$
@@
$\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{q}(Q_{R/2}(z))}\leqslant C\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{p}({Q_{R}(z))}}+C\|\hskip-6.4pt{{-}}g\|\hskip-6.4pt{{-}}_{L^{q}(Q_{R}(z))},$ (2.23)
@@
$\iint_{Q_{R}}\Delta u~\bar{u}\varphi^{10}\phi^{2}=\iint_{Q_{R}}|\nabla u|^{2}\varphi^{10}\phi^{2}+10\iint_{Q_{R}}\bar{u}\varphi^{9}\phi^{2}~\nabla u\cdot\nabla\varphi.$
@@
${\leqslant}$ $\delta J_{1}+C\delta^{-\frac{1}{2}}R^{-2}\left(\iint_{Q_{R}}|\nabla u|^{\frac{4}{3}}\right)^{\frac{3}{2}}.$
@@
${\leqslant}$ $\frac{C}{R}\|b\|_{L^{\infty}_{t}L^{2,\infty}_{x}(Q_{R})}\int_{-R^{2}}^{0}\|\bar{u}(t)\varphi^{\frac{9}{2}}\phi(t)\|_{L^{2}_{x}}\,\left\|\nabla\left(\bar{u}(t)\varphi^{\frac{9}{2}}\phi(t)\right)\right\|_{L^{2}_{x}}\,\text{\rm{d}}t$
@@
${\leqslant}$ $C\varepsilon^{-1}\delta J_{1}+C\varepsilon^{-1}\delta^{-\frac{1}{2}}R^{-2}\left(\iint_{Q_{R}}|\nabla u|^{\frac{4}{3}}\right)^{\frac{3}{2}}+\frac{\varepsilon}{2}\iint_{Q_{R}}|\nabla u|^{2}.$
@@
Let $d=2$. There exists a universal constant $\gamma>2$ only depending on $\|b\|_{L^{\infty}_{t}L^{2,\infty}_{x}}$ such that for any weak solution to (2.11) on $0,T]\times{\mathbb{R}}^{d}$ with $c=0$, it holds
@@
$\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{p}(Q_{{R}/{2}})}\leqslant\frac{1}{2}\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{p}(Q_{R})}+C_{6}\|\hskip-6.4pt{{-}}v\|\hskip-6.4pt{{-}}_{L^{1}(Q_{R})}+C_{6}\|\hskip-6.4pt{{-}}g\|\hskip-6.4pt{{-}}_{L^{p}(Q_{R})},$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]